The length of a rectangle is 5 cm greater than its width. The perimeter is 58 cm. Find the dimensions of the rectangle.
Answer by edjones(7311)

l=w+5
p=2(l+w)
58=2(w+5+w)=2(2w+5)=4w+10
4w+10=58
subtract 10 from both sides. 4w=48
divide 4 by both sides. w=12cm l=w+5=17cm
EdJones